OpenStudy (hippoharry):

You have some money to invest in one of two accounts. The first account pays 5% simple interest, and the second pays 4% compound interest. How would you decide which account to use?

OpenStudy (princeharryyy):

suppose you have $ 10 to invest

OpenStudy (hippoharry):

@princeharryyy more explaination please

OpenStudy (princeharryyy):

for year 1 si at 5% = 0.05*10 = 0.5 for 1 year ci at 4% = 0.04*10 = 0.4

OpenStudy (princeharryyy):

so if u have to invest for a year choice with si 5% is appropriate

OpenStudy (princeharryyy):

but,

OpenStudy (princeharryyy):

if you are planning for investing for more than 1 year we might have to reconsider

OpenStudy (hippoharry):

that doesnt really make any sense

OpenStudy (princeharryyy):

let us see what happens when we have two years. si for 2 years = 0.05*10*2 = 1 ci for 2 years = 0.04*10*1.04 + 0.04 = 0.816

OpenStudy (princeharryyy):

so even for two years you should choose si with 5 %

OpenStudy (a_burning_masquerade):

I don't know. I'm sorry. Percentages are my worst enemy.

OpenStudy (princeharryyy):

si for 3 years = 0.05*10*2 = 1.5 ci for 2 years = 0.04*10*1.04 + 0.04 + 0.04*10*1.0481.04= 1.2528 still for 3 years you should choose si with 5% interest on Si uptill 3 years is greater than Ci

OpenStudy (hippoharry):

what is the si and ci?

OpenStudy (princeharryyy):

So, if you continue going further in the same manner. there will be some year in which CI would be greater than Si. So until before that year you should invest in SI and as soon as that year arrives you start investing in Ci. Means it all depends on the number of years for which you have to invest.

OpenStudy (princeharryyy):

Si is simple interest and CI is compound interest

OpenStudy (hippoharry):

im sorry but can you explain it like your talking to a 5 year old cuz it still doesnt make sense, or either im really stupid

OpenStudy (princeharryyy):

si is a linear quantity while ci is exponential. So at some year say n there would come a time when CI would be greater than SI. at that point if you are investing up till year n-1 you should choose one with SImple Interest at 5% of if years are greater than or equal to that year n. you should invest in CI with 4% interest rate.

OpenStudy (princeharryyy):

I hope u understand what I am saying. If u don't Just copy this your teacher would know. Else, I can't help. Sorry about that.

OpenStudy (mathstudent55):

You shake your piggy bank at home and there is money in it. Your mommy tells you that you can take your money to a bank and start your very own account. You break your piggy bank, take all your money, and you put it in your pink Hello Kitty wallet. You jump into your mommy's car and your mommy fastens your seat belt safely. When you get to the bank, you are greeted by a nice lady who gives you a lollipop. Your mommy tells her that you are in the bank to open your first bank account. The nice lady tells you and your mommy that they have two different accounts for you. You can open an account that gives 5% simple interest, or you can open an account that gives 4% compound interest. You are confused, but your mommy tells you not to worry because your daddy is a smart accountant who knows all about interest rates and bank accounts. Your mommy calls your daddy and tells him what the bank lady said. Your daddy explains: Simple interest is interest you earn on the principal amount. That means it is interest that is earned only on the amount you deposit. No matter how much interest you earn, the amount of interest you will earn each year is sill based on the first amount you deposited. Compound interest is interest you earn on the amount you deposit and on the amount of interest you have already earned. As you earn more and more interest, you are earning interest on more money because with compound interest you earn interest on the interest. At the same interest rate, compound interest will always give you more interest. At a lower interest rate, compound interest may eventually give you more interest if you leave the money in the bank for long enough. The decision of which account to start depends on how long you plan to leave your money in the bank. By now, you are half asleep, and all you want is to either go back in the car, or find the nice lady again and ask her for another lollipop.